OpenStudy (anonymous):

A regulation basketball has a diameter of 9.4 inches. What is the volume of the ball? A) 19.69 in3 B) 92.53 in3 C) 434.89 in3 D) 3479.14 in3

OpenStudy (anonymous):

r = 9.4/2 = 4.7 V = 4/3 π r^3 V = 4/3 π (4.7^3) = 434.89 in3 The answer is C